OpenStudy (anonymous):

whats the numertator of the following rational expression a/b + 6/b = ?/b whats the numertator of the following rational expression a/b + 6/b = ?/b @Mathematics

OpenStudy (anonymous):

oh cmon man... a+6

OpenStudy (anonymous):

try and think a little bit

OpenStudy (anonymous):

its not too difficult

OpenStudy (anonymous):

thats not right though bruh thats why i asked

OpenStudy (anonymous):

on the homeowrk it says it wrong?

OpenStudy (anonymous):

homework

OpenStudy (anonymous):

well they are wrong bcuz that is right

OpenStudy (anonymous):

so thats why i keep asking

OpenStudy (anonymous):

the question is like this right?\[\frac{a}{b}+\frac{6}{b}=\frac{?}{b}\]

OpenStudy (anonymous):

a+6 is correct what does the hw say

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is definitely:\[a+6\]for an answer then.

OpenStudy (anonymous):

a+6 or 6+a. Is it automated homework (i.e. on a computer)? It may be expecting very specific input, such as 'a + 6' or '6 + a' or '6+a' or 'a+6' - it can get finicky sometimes.

OpenStudy (anonymous):

if you can, take a screen shot of the question, assuming its something on your computer.

OpenStudy (anonymous):

yea when doing math on a comp things tend to get weird when not putting in the way the computer wants it

OpenStudy (anonymous):

Which sucks. For example a*6 = a6 = 6a = (6)(a) = (6)a = a(6) = (a)6 etc etc. But the computer might only accept '6a' and nothing else, whereas a human would recognise it as correct.

OpenStudy (anonymous):

it must be -6